OpenStudy (booksforlife):

I could really use some help. 81^x=27^(x+2) I know the base is 3.

OpenStudy (saiyansparta):

\[(81)^x = (3^4)^x=3^{4x}\] and \[ 27^{(x+2)}=(3^3)^{(x+2)}=3^{3(x+2)}.\] Equating the two gives \[ 3^{4x}=3^{3(x+2)}\implies 4x=3(x+2)\implies 4x = 3x+6 \implies \boxed{x = 6}\]

OpenStudy (booksforlife):

Thank you so much @SaiyanSparta

OpenStudy (booksforlife):

Could you help me with one more?

OpenStudy (saiyansparta):

Go ahead and post it. :)